# UniCode: A Framework for Generating High-Quality Competitive Coding Problems

### v1

@@ 2.2.2 Output Generation
Establishing ground-truth outputs for novel problems is challenging. We devise a multi-stage pipeline (Figure 2) that mirrors rigorous human validation practices, separating inputs into small-scale and large-scale sets to balance reliability with computational feasibility.
Stage 1: Brute-Force Validation and Solver Filtration.
For small-scale inputs ($I_{s}$), where correctness can be verified tractably, we first generate a brute-force solver ($B$) via an LLM prompt. This solver, while inefficient, is optimized for correctness. We execute $B$ on all inputs in $I_{s}$ to create a trusted set of ground-truth pairs, forming our initial stress test suite. [^1] To mitigate the risk of a single faulty brute-force solver, we generate multiple candidates and use their consensus output, further enhancing reliability.
Next, we prompt various LLMs to produce a diverse set of $M$ optimized candidate solutions, $\{C_{1},\dots,C_{M}\}$. Each candidate $C_{j}$ is then validated against our stress test suite. A solution is admitted to our trusted solver pool $\mathcal{P}$ only if it correctly reproduces the brute-force output for every small-scale input:
$\mathcal{P}=\{C_{j}\mid C_{j}(i)=B(i)\text{ for all }i\in I_{s}\}.$
This filtration step ensures that only solutions proven correct on a baseline set of inputs are used for subsequent stages, yielding a pool of $N=|\mathcal{P}|$ efficient and pre-vetted solvers.
Stage 2: Consensus Validation on Large-Scale Inputs.
For large-scale inputs ($I_{\ell}$), which include all test cases generated by random, adversarial, and LLM-based synthesis methods, executing a brute-force solver is computationally infeasible (see Appendix appendix˜C). Instead, we leverage the trusted pool $\mathcal{P}$ of efficient solvers. For each input $i\in I_{\ell}$, we execute all $N$ solvers in $\mathcal{P}$ and collect their outputs. The final ground-truth output is determined by a strict majority vote: an output $o$ is accepted if it is produced by more than $\lfloor N/2\rfloor$ solvers.
Stage 3: LLM Adjudication for Disagreements.
If no single output achieves a strict majority, the input is flagged for adjudication. In this case, we identify the top two most frequent outputs, $o_{1}$ and $o_{2}$, and present them to a powerful LLM (e.g., o4-mini). The LLM is tasked with analyzing the problem statement, the input, and the conflicting candidate solutions to determine which output is correct. If the LLM provides a decisive judgment, its chosen output is accepted as the ground truth. If the conflict cannot be resolved, the input is discarded from the final test set to maintain its integrity. This final step allows us to resolve complex edge cases that might stump a purely algorithmic consensus. We validate each component in this pipeline can improve the accuracy of test case in Table˜1.
3 Benchmark Curation
This section describes how we construct the benchmark produced by our UniCode framework and how we validate the resulting problems and test suites. We first detail the data sources and curation pipeline used to assemble the benchmark, then report a human study assessing problem quality and clarity.
3.1 Benchmark structure
Seed problems.
We start from the TACO dataset, which aggregates over $25k$ programming problems with human-written solutions from platforms e.g., Codeforces, LeetCode, and CodeChef.
Filtering for competitive difficulty.
We filter the dataset by: (i) removing non-competitive problems; (ii) eliminating duplicates and problems with unclear input/output specifications.
Target coverage and tag taxonomy.
From the filtered pool, we target the 15 most prevalent algorithmic tags to ensure breadth while keeping consistent difficulty. Our tag set includes hierarchical refinements (e.g., graph $\rightarrow$ shortest-paths, flow; dp $\rightarrow$ knapsack, interval). The full taxonomy and mapping rules appear in Appendix D. The tags are assigned by an LLM, which we prompt to identify the 1-3 most relevant skills for each problem (see Appendix H).
Problem generation and selection.
We adopt the o4-mini-medium to generate a set of candidate problems based on our initial problem pool, using the UniCode methods described in §2. Each generated problem is paired with a test suite averaging 50 cases. To ensure the problems are challenging, we filter out any that can be solved perfectly by all our baseline models. Concretely, we run a panel of baselines as in  Table˜2 if every model achieves a $100\%$ pass rate on the test suite, the problem is removed. Our final benchmark contains 492 problems spanning the 15 target tags.
Test suites and constraints.
@@ 3.1 Benchmark structure
Each problem includes five components: description $D$, tag set $\mathcal{T}$, time limit $TL$, memory limit $ML$, and test cases $M$. Test cases are assembled as in §2.2 with a fixed composition of $|I_{\mathrm{rand}}|=20$, $|I_{\mathrm{adv}}|=20$, and $|I_{\mathrm{llm}}|=10$ (total $|M|=50$). The time limit $TL$ and memory limit $ML$ are determined by running validated, optimized solutions $\mathcal{O}_{\text{valid}}$:
@@
where $o^{\star}\;=\;\arg\min_{o\in\mathcal{O}_{\text{valid}}}T(o).$ We set safety factor $k=3$, and execute $T(\cdot)$ and $\mathrm{Mem}(\cdot)$ within a secure sandbox environment (Bytedance-seed et al., 2025). To facilitate analysis, we also store per-case provenance (generator type, random seed, and verifier decisions).
3.2 Quality of Generated Problems
Human evaluation protocol.
We assess clarity and solvability through a blinded rating study (see Appendix E). Five experienced annotators (competitive programmers/algorithm engineers, $\geq$5 years experience) independently rate a uniformly random sample of $50$ problems from the $492$-problem benchmark. We provide each rater with the problem statement and example I/Os. The primary metric is binary solvability (solvable / unsolvable): whether the statement is unambiguous and admits a well-defined solution within the stated constraints.
Results.
Across the 50 evaluated problems, the solvability rate is 98.0%. The average inter-annotator agreement rate with human assessments is 92.3%, indicating that the problem statements were clear and well-posed for expert readers.
@@ Results.
We will release problem statements, test suites (with provenance), and metadata (tags, generators, and prompts) to support reproducibility and downstream analysis.
3.3 Evaluating Test Case Quality
We evaluate test suites using two metrics: correctness (accepting valid solutions) and coverage (rejecting invalid ones). A robust test suite must excel at both. Let $S_{\text{correct}}$ and $S_{\text{incorrect}}$ be sets of correct and incorrect submissions. A test suite $M$ passes a submission $s$ if $s$ succeeds on all test cases $m_{i}\in M$, denoted $\text{pass}(s,M)=1$. We define:
$\text{Corr@N}=\frac{|\{\,s\in S_{\text{correct}}\mid\text{pass}(s,M)=1\,\}|}{|S_{\text{correct}}|},\quad\text{Cov@N}=\frac{|\{\,s\in S_{\text{incorrect}}\mid\text{pass}(s,M)=0\,\}|}{|S_{\text{incorrect}}|}.$
Table 1: Quality of generated test suites. We compare our full pipeline against the rStar-Coder baseline and ablations of our own method. Stage 1 (Brute-Force) uses brute-force solvers on small random inputs. MajVote (Unfiltered Solvers) applies a majority vote to all generated solutions without pre-filtering. MajVote (Filtered Solvers) uses only solutions that passed the Stage 1 stress test. Our Full Pipeline integrates all stages and input types. “Validated suites” is the percentage of problems for which a suite passed all validation checks.
Method Input types Correctness Coverage
Problems with
validated suites (%)
rStar-Coder (Liu et al., 2025) MajVote (Unfiltered) RandRand 86.9% 80.2% 94.3%
Ours Stage 1 (Brute-Force) $Rand_{s}$ 91.9% 81.5% 98.2%
MajVote (Unfiltered) $Rand_{l}$ & AdvAdv 86.7% 85.2% 93.9%
MajVote (Filtered) $Rand_{l}$ & AdvAdv 93.8% 84.3% 92.8%
Full Pipeline RandRand & AdvAdv 94.5% 86.0% 94.8%
Setup.
We test on 80 problems from the Test-Eval dataset (Yang et al., 2025), using suites of size $N=50$. Note that our approach to majority voting differs from rStar-Coder in its granularity. While rStar-Coder aggregates at the solution level and discards a problem unless a majority of solutions exhibit identical input–output behavior; our aggregation is performed per test case, so disagreement on individual cases does not invalidate the entire problem.
Results.
Our full pipeline significantly outperforms the baseline, achieving higher scores in both correctness (94.5% vs. 86.9%) and coverage (86.0% vs. 80.2%). We attribute this improvement to several key factors. Our Stage 1 filtering boosts correctness from 86.7% to 93.8% by using brute-force solvers on small inputs to eliminate flawed solutions. At the same time, we improve coverage by using adversarial inputs to uncover edge-case failures that are often missed by random inputs. Although our system does not achieve 100% accuracy, the mathematical proof in Appendix B demonstrates that the benchmark’s reported accuracy is still trustworthy. Finally, the validated-suite rate slightly exceeds the baseline because our method of per-test-case adjudication helps us retain reliable, yet difficult problems that would otherwise be discarded.
4 Evaluating LLMs on Algorithmic Problems
This section presents the experimental setup and results of our evaluation of large language models (LLMs) on the UniCode benchmark. We analyze model performance across various difficulty levels and compare the capabilities of different model series, providing insights into their strengths and weaknesses on code generation tasks.
4.1 Experimental Setup
Models
We evaluated 19 diverse large language models, including open-source and closed-source models with varied architectures and sizes. Our selection includes models optimized for reasoning, instruction-following, and code generation. Key models are GPT series (e.g., gpt-4.1-mini, gpt-4o), Gemini series (gemini-2.5-pro, gemini-2.5-flash), Claude series (claude-3.7-sonnet, claude-sonnet-4). Other competitive models from the DeepSeek and Qwen series, as well as llama-4-maverick and gemma-3-27b, are also included for a comprehensive comparison. We use a unified prompt with a model temperature of 0.2 across all models.
Model Easy Medium Hard Overall RandPass AdvPass $\Delta(A-R)$
o4-mini (high)* 94.9% 78.2% 21.6% 70.3% 80.1% 72.7% -7.4%
gpt-5 (medium)* 89.5% 77.6% 18.8% 67.7% 75.9% 70.0% -5.9%
o4-mini (medium)* 89.2% 73.6% 20.3% 66.1% 80.4% 67.5% -12.9%
google/gemini-2.5-pro* 94.0% 53.1% 8.5% 61.6% 74.7% 64.3% -10.4%
deepseek-v3.1 89.2% 59.8% 11.5% 60.5% 70.3% 64.4% -5.9%
deepseek-r1 80.3% 36.4% 5.1% 55.6% 69.6% 59.6% -10.0%
o3-mini (medium)* 86.2% 50.0% 6.0% 55.1% 66.8% 57.5% -9.3%
qwen3-235b-a22b 80.2% 39.7% 5.1% 53.5% 68.2% 56.0% -12.2%
gemini-2.5-flash* 81.4% 22.6% 4.8% 47.7% 61.8% 50.4% -11.4%
grok-3-mini* 77.8% 21.7% 3.3% 46.4% 61.8% 48.3% -13.5%
claude-3.7-sonnet* 76.2% 24.1% 2.4% 45.5% 63.9% 44.5% -19.4%
deepseek-chat-v3.1 82.7% 29.3% 3.9% 49.8% 61.0% 52.0% -9.0%
gpt-4.1-mini* 73.7% 20.9% 3.8% 42.4% 58.1% 45.4% -12.7%
gpt-4.1* 62.1% 21.8% 1.4% 36.5% 51.8% 39.4% -12.4%
qwen3-coder 66.5% 9.3% 0.0% 35.4% 53.9% 36.9% -17.0%
claude-sonnet-4* 60.7% 14.0% 2.0% 32.4% 48.3% 34.1% -14.2%
llama-4-maverick 51.3% 8.6% 0.0% 26.2% 41.3% 27.5% -13.8%
gpt-4o* 31.3% 2.2% 0.0% 15.4% 23.9% 15.2% -8.7%
qwen-2.5-32b-coder 27.2% 2.2% 0.0% 13.4% 26.8% 11.5% -15.3%
gemma-3-27b-it 26.1% 2.2% 0.0% 13.1% 21.2% 14.0% -7.2%
llama-3.3-8b-instruct 11.2% 1.1% 0.0% 5.5% 12.8% 5.8% -7.0%
Table 2: Pass@1 performance on the UniCode benchmark. The table is split by reasoning (upper) and non-reasoning (lower) models. Closed models are marked with *. Models are ranked by overall average Pass@1 in descending order. RandPass and AdvPass denote pass rates on random and adversarial test cases respectively. $\Delta(A\!-\!R)$ denotes the gap between AdvPass and RandPass.
4.2 Pass@1 performance overview
We report pass@1 results on the UniCode benchmark, which clearly ranks models by performance and difficulty levels while exposing their brittle failure modes. Our key findings are:
UniCode is challenging and highly discriminative.
Even the top model, o4-mini-high, attains an overall pass@1 of 70.3%, while the weakest model (llama-3.3-8b-instruct) reaches only 5.5%, a significant gap of 64.8% points. Performance collapses sharply on hard problems: several models, such as gpt-4o and llama-4-maverick, record 0.0% on the hard split, underscoring the benchmark’s capacity to distinguish multi-step algorithmic competence. The results further reveal that reasoning-oriented models consistently outperform others. The top six reasoning models achieve an average pass@1 of 57.3%, which is more than double the 30.0% average of the non-reasoning models. This substantial gap highlights the value of architectures and fine-tuning specifically for structured, multi-step reasoning in competitive coding tasks. The ranking results from the UniCode benchmark show over 90% alignment with other popular code benchmarks, as detailed in Appendix A.3
Models consistently struggle with challenging test cases.
We contrast performance on randomly sampled inputs (RandPass) with adversarially generated inputs (AdvPass). Across all models, the average drop from RandPass to AdvPass is 11.2% points, indicating widespread brittleness to edge cases. The most significant declines occur in claude-3.7-sonnet with a 19.4% point decrease and qwen3-coder with a 17.0% point decrease. Interestingly, while o4-mini-medium achieves a RandPass rate similar to o4-mini-high, it exhibits a substantial 5.2% point gap in AdvPass at 67.5% versus 72.7%. This disparity suggests that higher-capacity reasoning models possess stronger abilities to handle complex corner cases, thereby improving adversarial robustness.
Competitive open-source models are on the rise.
While closed-source models generally lead in overall performance, a select few open-source models are proving to be serious contenders. Models like DeepSeek-v3.1 and Qwen3-235b-a22b demonstrate robust capabilities, showcasing the immense potential of the open-source community. Notably, DeepSeek-chat-v3.1, a non-reasoning model, achieves a overall pass@1 of 49.8%, outperforming several reasoning-oriented models like gemini-2.5-flash. The strong performance of these open-source models provides immense value to the research community. They offer transparency and scientific understanding, which allows researchers to conduct deeper investigations into model architecture, training data, and limitations, fostering innovation and accelerating progress in competitive programming.
We further explore the pass@k performance in Appendix A.2. The performance for o4-mini-medium rises from $66.1\%(k=1)$ to $83.5\%(k=10)$ but plateaus after $k=7$, indicating diminishing returns. The final pass rate underscores the benchmark’s challenge even with extensive sampling.
4.3 Evaluating Generalization in LLMs
Figure 3: Performance comparison across question sets. SeedQS denotes the original seed questions, ShadowQS represents the semantically equivalent shadow questions generated from SeedQS, and CodeGenQS refers to the novel question set created by our UniCode method.
This section we evaluate whether models can flexibly apply learned knowledge to solve novel problems, rather than simply retrieving memorized solutions. We construct three problem sets derived from 50 problems in LiveCodeBench v1  [^2] Initial Release: May 2023 - March 2024., an early version with potential data contamination risk. Seed questions: the original 20 problems. Shadow questions: reformulated versions with altered wording, added stories, or distracting details, while preserving the same logical core. UniCode questions: new problems generated by UniCode (section˜2.1), requiring transfer and flexible use of seed knowledge. Each dataset contains 48 problems. We evaluated pass@1 performance across four LLMs, yielding key insights into their generalization abilities:
Robustness to Superficial Changes Most LLMs perform nearly identically on seed and shadow questions. This suggests a robust capacity to generalize across variations in problem descriptions. Models effectively distill the core logical structure, remaining unaffected by superficial details or narrative changes, indicating a reliance on deeper comprehension over shallow pattern matching.
Algorithmic Generalization Remains a Challenge All LLMs exhibited a significant performance drop on UniCode questions. The decline was particularly pronounced in claude-3.5-sonnet (0.70 → 0.21) and gpt-4.1-mini (0.83 → 0.40). This indicates that while models handle narrative variations well, they struggle when faced with problems requiring novel or composite algorithmic reasoning. Qwen3-235b-a22b and o4-mini-medium demonstrated relatively stronger generalization, suggesting a more flexible problem-solving architecture.
4.4 Performance Across Algorithmic Paradigms

### v2

@@ 3.2 Ground-Truth Construction
Establishing ground-truth outputs for novel problems is challenging. We devise a multi-stage pipeline (Figure 2) that mirrors a rigorous human validation process.
Stage 1: Brute-Force & Solver Filtration We generate a brute-force solver $B$ via LLM to create ground-truth pairs for small-scale inputs ($I_{s}$). Multiple candidates and consensus ensure $B$’s reliability. We then prompt LLMs for $M$ optimized candidate solutions $\{C_{1},\dots,C_{M}\}$. A candidate enters the trusted pool $\mathcal{P}$ only if it matches $B$ on all $I_{s}$:
$\mathcal{P}=\{C_{j}\mid C_{j}(i)=B(i)\;\forall i\in I_{s}\}.$
Stage 2: Consensus on Large-Scale Inputs For large-scale inputs ($I_{\ell}$) where brute-force is infeasible, we use the pool $\mathcal{P}$. The ground truth for $i\in I_{\ell}$ is determined by a strict majority vote ($\lfloor N/2\rfloor$) among the $N$ optimized solvers.
Stage 3: LLM Adjudication If no majority exists, the top two outputs ($o_{1},o_{2}$) are sent to a high-reasoning LLM (e.g., o4-mini) for analysis. If the LLM yields a decisive judgment, that output is accepted; otherwise, the input is discarded to ensure data integrity. We validate that each component improves test case accuracy in Table 3.
Figure 3: Model performance across reasoning variants. Red dashed lines denote seed problem performance. Key observations: (1) All models exhibit performance variance (e.g., claude-sonnet-4 shows a gap of $60\%$), suggesting substantial disparities in reasoning capabilities across different dimensions. (2) Even reasoning-optimized models (o4-mini, deepseek-r1) are vulnerable to sequential integration, concept fusion and efficiency scaling, which exposes a critical fragility of multi-step reasoning under novel constraints.
4 Benchmark Curation and Leaderboard
This section describes the construction and validation of the UniCode benchmark. We first present our data curation pipeline and a human study evaluating problem quality, followed by a comprehensive leaderboard overview.
4.1 Data Pipeline and Quality
Problem Generation
We curated a dataset of $25,000$ seed problems aggregated from various platforms (e.g., LeetCode, CodeForces). To ensure competitive quality, we removed non-competitive tasks, duplicates, and unclear specifications. Tags were assigned hierarchically (e.g., graph $\rightarrow$ shortest-paths, flow) by an LLM identifying 1–3 core skills per problem (see App. A.6 and B.8). To minimize potential generator bias, as illustrated in App. B.3, we used o4-mini and deepseek-r1 to generate 492 candidate problems following the methods in §2. We excluded low-difficulty problems that were solved perfectly by all baseline models.
Test Suites and Constraints
@@ 4.1 Data Pipeline and Quality
Each problem includes five components: description $D$, tag set $\mathcal{T}$, time limit $TL$, memory limit $ML$, and test cases $\mathcal{C}$. The time limit $TL$ and memory limit $ML$ are determined by running validated, optimized solutions $\mathcal{O}_{\text{valid}}$:
@@
where $o^{\star}\;=\;\arg\min_{o\in\mathcal{O}_{\text{valid}}}T(o).$ We select the minimum runtime across validated reference solutions to avoid loose time limits and multiply it by a conservative safety factor $k=3$ to accommodate variations in alternative correct implementations. We then execute $T(\cdot)$ and $\mathrm{Mem}(\cdot)$ within a secure sandbox environment (Bytedance-seed et al., 2025).
Test Suite Quality
To evaluate the quality of the generated problem variants, we assessed 113 sampled problems via a blinded study with five expert annotators. The benchmark achieves a 98.2% validity rate and 92.3% inter-annotator agreement, confirming technical clarity (See App. B.4 for detail). To evaluate the test cases generated by the stress-driven pipeline, we assess correctness (valid solutions accepted) and coverage (invalid solutions rejected). Our ablation study in App. B.1, using Test-Eval (Yang et al., 2025b), demonstrates the effectiveness of our pipeline: reaching 94.5% correctness (vs. 86.9% baseline) and 86.0% coverage (vs. 80.2% baseline), as shown in Table 3. While automatic generation is not error-free, our mathematical proof in App. B.5 confirms that the generated benchmark remains statistically reliable for evaluation.
@@
We will release problem statements, test suites, and metadata (tags, generators, and prompts) to support reproducibility and downstream analysis.
4.2 UniCode Leaderboard
To offer a macroscopic perspective on the UniCode landscape, we evaluate 19 state-of-the-art LLMs across various architectures, parameter scales, and reasoning capabilities.
As shown in Table 1, UniCode is both highly challenging and discriminative, with overall pass@1 scores ranging from 70.3% (o4-mini-high) to 5.5% (llama-3.3-8b-instruct). Performance collapses on the hard split, where several models record 0.0%, underscoring the benchmark’s difficulty. Reasoning-oriented models lead the rankings, validating the effectiveness of test-time compute scaling for complex logical inference. Our results show over 90% alignment with uncontaminated benchmarks (App. A.4), confirming that UniCode is robust against data contamination and provides an unbiased assessment of model performance.
Models generally struggle more with adversarial and corner cases than random generation. For instance, qwen3-235b-a22b show pronounced sensitivity, with $\Delta_{\mathrm{Corn}}$ reaching $13.3\%$. Conversely, deepseek-v3.1 (thinking) exhibits superior stability with minimal performance drops. Furthermore, cost–performance analysis identifies o4-mini (high) as highly efficient, achieving top-tier pass@1 at \$$0.0269$ per problem - nearly $7.5\times$ more cost-effective than gemini-2.5-pro for similar accuracy. While these results reveal UniCode’s difficulty, they do not pinpoint specific failure modes. The following section decomposes performance across five reasoning axes to uncover where and why models fail.
5 In-Depth Code Reasoning Analysis
In this section, we investigate the fragility of LLM code reasoning across diverse variants and introduce a fine-grained taxonomy to categorize the origins of these failures.
5.1 How Fragile is Code Reasoning in LLMs?
To better understand the reasoning ability of LLMs, we curated 132 new problems using livecodebench v1  [^1] Initial Release: May 2023 – March 2024. seed tasks, where models previously excelled. We evaluated a selection of reasoning-specialized, general-purpose, and open-source models across various performance tiers.
General Performance Drop and High Variance
As shown in Figure 3, all models suffered significant performance declines, revealing their fragility when handling structural and conceptual shifts in reasoning tasks. We observe that even state-of-the-art LLMs fail to achieve comprehensive mastery across diverse reasoning paradigms. Furthermore, models exhibit non-negligible variance across test sets; for instance, claude-3.5-sonnet shows a performance gap exceeding 60% between scenarios. This imbalance suggests that aggregate scores often mask significant deficiencies in generalization.
Failure Under Structural Reasoning Alterations
While some LLMs remain robust against narrative perturbations, all models suffer sharp performance declines when the underlying reasoning graph topology is altered. Constraint modifications consistently degrade performance, highlighting the difficulty of preserving logical coherence under novel requirements. Efficiency scaling is particularly challenging: even the specialized o4-mini experiences a 37% drop, marking its weakest dimension. The most significant failures occur in sequential reasoning and concept fusion, indicating persistent difficulty in composing multiple logical components and maintaining long causal chains. Notably, gemma-3-27b-it scored a negligible 1.3% in sequential tasks, representing a near-total loss of functional capacity. Model performance exhibits a continuous declining trend as the augmentation depth increases (see App. A.5).
Seed-problem Regression Phenomenon.
We observe that models often default to original seed-problem logic rather than reasoning from updated task specifications (see Figure 5). This behavior suggests a reliance on heuristic shortcuts rather than rigorous logical deduction. For example, a model may employ a simple parity count because it recognizes a palindrome sub-task, yet fail to integrate new constraints. Alternatively, models rely on “low-efficiency templates” suitable for base problems without accounting for increased input scales. Future research should focus on enhancing models’ zero-shot adaptation to novel constraints.
Figure 4: Error distributions across LLMs. Modeling and complexity errors are the primary failure modes, revealing inherent algorithmic weaknesses. (See Figure 5 for the detailed case study.)
Figure 5: Case study of error patterns in code reasoning. Model failures are not stochastic but exhibit a seed-problem regression. Even when new problem invalidate original P1/P2 logic, LLMs frequently revert to memorized modeling paradigms or suboptimal complexities that were sufficient for the base problems. This suggests that LLMs often rely on heuristic shortcuts rather than true logical synthesis.
5.2 Where Do LLMs Fail in Code Reasoning?
To provide a more comprehensive overview of why LLMs fail, we introduce a systematic taxonomy to analyze their root causes. Modeling error: selecting an incorrect algorithmic paradigm for the task (e.g., opting for dynamic programming when a greedy approach suffices). Logic bug: implementation structural flaws (e.g., incorrect conditional or improper variable handling). Indexing/caching bug: incorrect cache sizing, array bounds violations, or failed boundary checks. Complexity error: utilizing suboptimal algorithms when constraints demand more efficient solutions. Others: minor implementation oversights, including output formatting errors or library usage. We prompt LLMs to categorize errors by providing failed code, passed solutions and failed test cases (see App. Section B.9).
Through a fine-grained analysis of failed cases, we observe an “overhead” phenomenon: models like gemma-3-27b-it and gpt-4o-mini rarely fail due to isolated mistakes. Instead, they exhibit cascading failures where multiple errors occur simultaneously. This indicates that when a task exceeds a certain complexity threshold, the model’s logical coherence may break down, resulting in repetitive, nonsensical code snippets or hallucinatory logic.
As shown in (Figure 4), indexing and others errors account for the smallest fraction of failures. Although dominant error types vary, modeling errors remain a primary challenge, indicating persistent difficulty in problem conceptualization and algorithm selection. Another major contributor is complexity error: models such as gemini-2.5-flash and gemma-3-27b-it struggle with time complexity analysis and appear insensitive to resource constraints. These results suggest that while LLMs are proficient at syntax, they lack a robust grasp of algorithmic efficiency and conceptual modeling.
6 Related Work
Competitive Coding
LLM code generation evaluation is a rapidly evolving field (Jaech et al., 2024; Li et al., 2023; Guo et al., 2024; Hui et al., 2024; Zhang et al., 2023; Guo et al., 2025; Li et al., 2022b; Shao et al., 2024; Allal et al., 2023; Zhao et al., 2025), as code is increasingly viewed as a potential source of reasoning ability (Fu et al., 2022; Li et al., 2022a). Foundational benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021) use standalone patterns, but their static nature risks data contamination and statistical shortcuts (Oren et al., 2023; Golchin and Surdeanu, 2023; Riddell et al., 2024; Roberts et al., 2023; Tang et al., 2024). Recent benchmarks (Li et al., ; Gu et al., ; Zhu et al., 2025; Chambon et al., ) introduce complex competitive problems to test algorithmic reasoning, yet updates remain slow and patterns fixed (Zheng et al., 2025b; Jain et al., 2024). We study open-ended tasks (Stanley and Lehman, 2015; Clune, 2019; Lehman and Stanley, 2011) explored in recent benchmarks (Xie et al., 2024; Heesch et al., 2025; Chen et al., 2025), which have limited difficulty and reasoning. UniCode generates infinite competitive programming problems via deep structural shifts to probe true code reasoning.
Automated Test Generation
Accurate algorithmic assessment requires scalable test generation, which often relies on manual curation (Chen et al., 2021; Hendrycks et al., 2021; Austin et al., 2021; Li et al., ; Quan et al., 2025). While some use LLMs for test synthesis (Chen et al., 2022; Schäfer et al., 2023; Liu et al., 2023a; Wang et al., 2025; Jain et al., 2024), they frequently depend on existing reference solutions (Schäfer et al., 2023; Tufano et al., 2022; Chen et al., 2022; Liu et al., 2023b). Other works (Chou et al., 2025; Lops et al., 2025; Anand et al., 2013; Sofokleous and Andreou, 2008; Tufano et al., 2020; Swain et al., 2012) generates problems from code but lacks algorithmic depth. UniCode introduces a reliable, stress-driven pipeline enabling robust evaluation for novel problems without ground-truth code.
Generative and Augmented Evaluation
To combat benchmark stagnation (Oren et al., 2023; Roberts et al., 2023), generative evaluation creates dynamic tasks, as seen in game environments (Zheng et al., 2025a), medical diagnosis (Lin et al., 2025), logical reasoning (Parmar et al., 2024), synthetic data structures (Zhu et al., 2023), and agentic tasks (Shi et al., 2025). We adapt this to algorithmic reasoning, using various perturbations and structural variations that may change the underlying reasoning graph (Pei et al., 2025; Huang et al., 2025; Wu et al., 2021) to expose reasoning fragility. Unlike traditional instruction-tuning or Self-Instruct (e.g.,  (Xu et al., 2024; Luo et al., 2023)), or recent robustness benchmarks like (Stolfo et al., 2023; Mirzadeh et al., 2024; Li et al., 2024a; Wang and Zhao, 2024; Yang et al., 2025a; Ramezanali et al., 2025; Orvalho and Kwiatkowska, 2025; Patel et al., 2024) which focus on surface-level perturbations or the linear extension of reasoning steps, UniCode targets underlying algorithmic transformations to disrupt memorized logic and evaluate true generalization.
7 Conclusion
In this paper, we introduce UniCode, a comprehensive and novel generative framework designed to address the “evaluation paradox” in current Large Language Model (LLM) coding benchmarks. To disrupt models’ reliance on statistical shortcuts, we implement multi-dimensional augmentations targeting structural, compositional, and conceptual shifts, supported by a stress-driven synthesis pipeline that ensures a scalable and contamination-resistant evaluation. Our findings reveal a significant 31.2% performance collapse across state-of-the-art LLMs, exposing a critical gap where high scores on static benchmarks fail to translate into robust logical adaptability. Ultimately, this research highlights the fragility of current models under novel logic patterns and underscores the urgent need for reasoning-oriented development in AI coding agents to bridge the gap between benchmark performance and real-world applicability.
Ethics Statement
This paper introduces UniCode, a framework designed to advance the field of machine learning by fostering genuine reasoning-oriented code intelligence. By exposing the “evaluation paradox” where models rely on statistical shortcuts, our work provides a critical foundation for developing more robust, generalizable AI agents. Beyond technical evaluation, UniCode contributes to the broader goal of building reliable and fair AI systems by systematically identifying logic fragilities. Ultimately, this research redefines benchmarking standards and facilitates the creation of safer AI technologies capable of handling complex, real-world reasoning tasks with higher fidelity and transparency.
Appendix A Extended Experimental Results and Diagnostics
A.1 UniCode Efficiency Leaderboard
To illustrate the performance-cost efficiency frontier and assist users in selecting models that balance budget with performance, Figure 6 presents the UniCode leaderboard. This plot maps the Pass@1 score against the average cost per problem for various models, including: gpt-5-2025-08-07, o4-mini-2025-04-16 high, o4-mini-2025-04-16 medium, gemini-2.5-pro, deepseek-v3.1-thinking, deepseek-r1-0528, o3-mini-2025-01-31, qwen3-235b-a22b, gemini-2.5-flash, grok-3-mini, claude-3.7-sonnet:thinking, deepseek-chat-v3.1, gpt-4.1-mini-2025-04-14, gpt-4.1-2025-04-14, qwen3-coder, claude-sonnet-4-20250514, llama-4-maverick:free, gpt-4o-2024-11-20, qwen-2.5-32b-coder, and llama-3.3-8b-instruct.
Figure 6: UniCode Leaderboard: Performance vs. cost efficiency across various models.
A.2 Performance Across Algorithmic Paradigms